# End-to-End Action Segmentation Transformer

### v1

@@ Abstract
Existing approaches to action segmentation use pre-computed frame features extracted by methods which have been trained on tasks that are different from action segmentation. Also, recent approaches typically use deep framewise representations that lack explicit modeling of action segments. To address these shortcomings, we introduce the first end-to-end solution to action segmentation – End-to-End Action Segmentation Transformer (EAST). Our key contributions include: (1) a simple and efficient adapter design for effective backbone fine-tuning; (2) a segmentation-by-detection framework for leveraging action proposals initially predicted over a coarsely downsampled video toward labeling of all frames; and (3) a new action-proposal based data augmentation for robust training. EAST achieves state-of-the-art performance on standard benchmarks, including GTEA, 50Salads, Breakfast, and Assembly-101. The model and corresponding code will be released.
@@ 1 Introduction
Action segmentation is a basic vision problem that involves labeling frames of an untrimmed video with their corresponding action classes.
The problem poses many challenges, including the high cost of framewise annotation, the inherent ambiguity of action boundaries,
and the significant computational demands of processing videos.
Recent approaches typically address these challenges by using pre-computed frame features, e.g., I3D [6] or TSM [22]. Such features are known to be suboptimal [8], because they have been previously extracted by other methods trained on vision tasks that are different from action segmentation. Furthermore, memory and computational constraints lead most recent approaches to focus on deep framewise representations [9, 38, 24] that lack explicit modeling of action instances (with few exceptions [26, 36] that increase complexity). Consequently, they neglect the bottom-up/top-down integration of frame and action representations, which was once essential in traditional frameworks [5, 28, 30, 1, 15]. Finally, as datasets for action segmentation are significantly smaller than for other tasks (e.g., action recognition), previous work resorts to data augmentation and self-supervised learning [2, 19, 27]. However, these methods typically augment only local frame features, and do not augment action instances.
@@
To address these shortcomings, we design End-to-End Action Segmentation Transformer (EAST) and make the following three key contributions, as illustrated in Fig. 1.
First, we enable efficient end-to-end training of EAST by introducing lightweight Contract-Expand Adapter (CEA) adapters into a large-scale backbone network. CEA applies feature compression and expansion around depth-wise convolutions to reduce complexity. This allows efficient fine-tuning of the backbone to extract multiscale frame features which are optimized for action segmentation, rather than for other vision tasks. To our knowledge, EAST is the first end-to-end action segmenter.
Second, unlike most recent approaches that refine framewise labeling in multiple stages without accounting for the temporal intervals of action instances, EAST performs action segmentation by detection. This offers two key advantages: efficiency – by detecting action proposals on a coarsely downsampled video instead of all frames – and improved framewise classification through explicit reasoning about detected action instances. As shown in Fig. 1, EAST consists of: (i) a detector that predicts action proposals over coarsely sampled frames; (ii) an aggregator that combines the proposals to predict a class distribution of every frame at the original (unsampled) frame-rate; and (iii) a refinement module for the final framewise labeling. By treating temporally sampled frames as action queries, EAST can efficiently and accurately predict action boundaries over the downsampled input. It is worth noting that temporal downsampling does not affect the ground truth or evaluation, as our boundary regression is specified relative to timestamps within the video. These boundary detections are mapped to the original full frame rate, and serve as useful constraints for final action segmentation over all frames.
Third, we introduce a new proposal-based data augmentation method to enhance EAST training. In training, our experiments show that the high-confidence action proposals generally align well with the ground truth, as intended. However, the domain gap between training and test data increases the uncertainty of action proposals generated by EAST on test videos. To improve robustness, we train EAST under higher-uncertainty conditions by conveniently integrating our proposal-based data augmentation in the end-to-end training. Our method randomly removes a subset of the most confident proposals and passes the remaining higher-uncertainty ones to the aggregator to learn a more reliable integration of action proposals for the final frame labeling. Unlike existing data-augmentation techniques that mask, shuffle, or manipulate video frames, our approach operates directly on action proposals.
@@ 3.1 Contract-Expand Adapter
$\bm{\bar{x}}$ $=\bm{\sigma}({\bm{W}_{\rm down}^{\rm\top}}\cdot\bm{x}),$ (1)
$\bm{\bar{x}}_{c}$ $=\text{contract}_{HW}(\bm{\bar{x}}),$
$\bm{\hat{x}}_{m}$ $={\bm{W}_{\rm mid}^{\rm\top}}\cdot\operatorname{\textbf{DWConv}}_{k}(\bm{\bar{x}}),$
$\bm{\bar{x}}_{e}$ $=\text{expand}_{HW}(\bm{\hat{x}}_{m}),$
$\bm{\hat{x}}$ $=\bm{\bar{x}}_{e}+\bm{\bar{x}},$
$\bm{x^{\prime}}$ $=\alpha\cdot{\bm{W}_{\rm up}^{\rm\top}}\cdot\bm{\hat{x}}+\bm{x},$
@@ 3.2 Low-Frame-Rate Action Detection
To enable efficient end-to-end training with temporal downsampling, we adopt a segmentation-by-detection framework, departing from SOTA approaches. Action proposals are predicted on coarsely sampled frames and then integrated at the original high frame rate for framewise classification. Inspired by anchor-free detectors (e.g., FCOS [33] and ActionFormer [40]), we treat each sampled frame as a query for its corresponding action proposal. This generates high-quality action proposals from low-FPS input. Compared to SOTA methods that rely on separate frame and action branches [26] or learnable queries [4], our approach significantly simplifies training by directly predicting action instances from sampled frames.
@@ 3.4 Training Loss Functions
where $\mathcal{L}_{\text{c}}$ denotes the focal loss [23]; $y_{i}$ is the ground-truth class; $\lambda_{\text{r}}$ is a weighting hyperparameter; $\mathbbm{1}(\cdot)$ is a binary indicator that equals 0 if $i$th frame is classified as background, and 1, otherwise; $T_{+}$ is the total number of sampled frames classified as an action $T_{+}=\sum_{i=1}^{T^{\prime}}\mathbbm{1}(\hat{y}_{i})$; $\mathcal{L}_{\text{r}}$ is the DIoU loss [41] for regression; $[t_{n(i)}^{s},t_{n(i)}^{e}]$ denotes the time interval of $n$th ground-truth action instance closest to the predicted interval $[\hat{t}_{i}^{s},\hat{t}_{i}^{e}]$ of $i$th action proposal. Note that the regression loss is applied only to action proposals classified as an action excluding background.
@@ 4 Results
Metrics. As in SOTA action segmentation approaches, we use framewise classification accuracy (Acc), edit score (Edit), and F1-scores (F1@$\{10,25,50\}$) at overlap thresholds of 10%, 25%, and 50% [18]. Edit score measures similarity between the predicted and ground-truth action sequences. F1-scores evaluate localization of action instances. We also report the average precision at Intersection over Union (IoU) thresholds of $\{0.3,0.4,0.5,0.6,0.7\}$ and the mean average precision (mAP) of our action proposals predicted by EAST’s detector.
@@ 4.1 Features and Training Efficiency
The SOTA methods use precomputed I3D features [6] for Breakfast and TSM features [22] for Assembly101. To ensure consistency in the frame features across the SOTA methods and EAST, we extracted MAEv2 frame features for all the methods, including ours, using the ViT-G backbone pretrained with VideoMAEv2 [35]. Tab.  1 shows that the SOTA methods improve performance on Breakfast when using precomputed MAEv2 features, while LTContext with MAEv2 features fails to do so on Assembly101. On both datasets, end-to-end trained EAST using the same pretrained ViT-G achieves superior performance.
@@
(’ICCV23) MAEv2 80.6 75.7 64.1 75.2 76.6
FACT [26] I3D 81.4 76.5 66.2 79.7 76.2
(’CVPR24) MAEv2 80.6 75.9 65.3 78.9 77.2
EAST MAEv2 85.3 81.2 71.4 83.3 81.7
Table 1: Impact of feature representation on Breakfast. SOTA methods use precomputed I3D features on Breakfast. To ensure a consistent feature representation, we compare SOTA and EAST using MAEv2 frame features extracted by the ViT-G backbone pretrained with VideoMAEv2 [35] across all the methods. Results are averaged across four standard splits.
@@
To evaluate training efficiency, in Tab. 3, we compare per-epoch training time, total training time, and GPU memory usage of the SOTA methods and EAST on Breakfast using NVIDIA H100 GPUs. Note that for this comparison the SOTA methods exclude the time required for frame feature extraction, whereas we include it, as it is an integral part of our approach. Tab. 3 shows that EAST achieves a total training time comparable to LTContext – the fastest feature-based Transformer model – despite directly processing RGB frames, while LTContext relies on precomputed features. Also, EAST’s training converges significantly faster in only 30 epochs, whereas LTContext requires five times more epochs. This highlights EAST’s efficiency in end-to-end training. Tab. 3 also shows that higher-performing models generally consume more memory, with EAST using approximately twice the memory of FACT. Considering that it requires 6GB of memory for the ViT-G backbone to extract per frame features, our memory usage is highly efficient compared to SOTA. This efficiency is due to our segmentation-by-detection framework which enables us to work with low frame-sampling rates at the input, unlike SOTA (see results in the supplement).
@@ 4.3 Aggregator and Sensitivity to FPS
Dataset Method F1@{10,25,50} Edit Acc
GTEA D 87.3 84.7 75.9 82.5 86.3
D+MSTCN [9] 93.2 93.0 88.7 92.1 83.8
D+ASFormer [38] 86.8 84.0 74.3 83.1 83.6
EAST 95.6 95.3 90.3 95.2 86.9
Breakfast D 76.5 72.2 63.4 72.3 78.4
D+MSTCN [9] 83.7 78.7 68.7 81.4 78.6
D+ASFormer [38] 83.9 79.5 68.7 81.6 78.8
EAST 85.3 81.2 71.4 83.3 81.7
Table 7: Action segmentation results using different refinement networks. D: Segmentation results obtained by aggregating action proposals as in (2) without any refinement. D+MSTCN and D+ASFormer: Results after refinement using MS-TCN and ASFormer, respectively.
@@ 4.4 Comparison with SOTA
ViT-G fine-tuned EAST 92.4 91.2 87.6 87.5 91.2
@@
ViT-G fine-tuned EAST 85.3 81.2 71.4 83.3 81.7
@@ 4.5 Qualitative Results
Fig. 4 compares of EAST segmentation results with the ground truth and SOTA on sample vides from GTEA, Breakfast, and Assembly101. In the top example video from GTEA, EAST successfully detects all actions, whereas the other methods miss at least one action. On the middle example video from Breakfast, the SOTA methods exhibit varying degrees of missed actions, spurious predictions, or oversegmentation. In contrast, EAST provides high-quality segmentation closely aligned with the ground truth. Finally, in the bottom video from Assembly101, EAST fails to detect a brief action instance and incorrectly identifies an action boundary too early – both of these are ambiguous edge cases that are challenging to discern even through visual inspection.
Figure 4: Segmentation results on sample videos from GTEA (top), Breakfast (middle), and Assembly101 (bottom). For each video sequence, the top row shows the color-coded ground-truth action instances, middle row shows EAST output, and bottom row visualizes SOTA results generated using publicly available models.
@@ 5 Conclusion
We have introduced EAST – the first fully end-to-end trainable action segmenter. EAST performs segmentation by detection, which enables temporal downsampling of input videos, significantly reducing computational costs. EAST takes RGB frames, sampled at a low frame rate, as input to a large-scale backbone. The backbone is fine-tuned using our Contract-Expand Adapter (CAE). CAE is especially effective in reducing the computational costs of end-to-end training by leveraging spatial pooling. The backbone features are passed to the detector to predict action proposals, which are then aggregated and refined to produce final framewise labeling at the original unsampled frame rate. We have also specified a novel proposal-based data augmentation that increases uncertainty of the detector’s output during training, effectively simulating test-time conditions. EAST outperforms prior work across all metrics on the GTEA, 50Salads, Breakfast, and Assembly101 datasets, while maintaining comparable processing times, even with the additional step of extracting frame features by the backbone. We have conducted a comprehensive ablation study to evaluate EAST’s performance under various configurations. Additional results are presented in the supplemental material. Code and models will be released.
@@ 6 Additional Results
SOTA Methods with Low FPS Input. The impact of video downsampling on the SOTA methods and EAST is evaluated on Breakfast in Table 14. The SOTA methods are trained using I3D frame features sampled at 1 FPS, with their output framewise classification subsequently upscaled to the original 15 FPS, as per their reported evaluation setting. Table 14 shows significant performance degradation for all the SOTA methods when working with the low FPS at the input. As shown in Tables 14 and 8, EAST maintains the best performance at low FPS rates and improves as the frame rate increases, subject to memory and compute constraints.

### v3

@@ Abstract
Most recent work on action segmentation relies on pre-computed frame features from models trained on other tasks and typically focuses on framewise encoding and labeling without explicitly modeling action segments. To overcome these limitations, we introduce the End-to-End Action Segmentation Transformer (EAST), which processes raw video frames directly – eliminating the need for pre-extracted features and enabling true end-to-end training. Our contributions are as follows: (1) a lightweight adapter design for effective fine-tuning of large backbones; (2) an efficient segmentation-by-detection framework for leveraging action proposals predicted over a coarsely downsampled video; and (3) a novel action-proposal-based data augmentation strategy. EAST achieves SOTA performance on standard benchmarks, including GTEA, 50Salads, Breakfast, and Assembly-101.
@@ 1 Introduction
Action segmentation is a basic vision problem that involves labeling frames of an untrimmed video with their corresponding action classes. The problem poses many challenges, including the inherent ambiguity of action boundaries, and significant computational demands of processing long video sequences.
Recent approaches typically address these challenges by using pre-computed frame features, e.g., I3D [6] or TSM [23]. Such features are known to be suboptimal [8], because they have been previously extracted by other methods trained on vision tasks that are different from action segmentation (e.g., action recognition). Furthermore, due to memory and computational constraints, most approaches focus on framewise representations [9, 39, 25] that lack explicit modeling of action instances (with few exceptions [27, 37] that increase complexity). Consequently, they neglect the bottom-up/top-down integration of frame and action representations, which was once essential in traditional frameworks [5, 29, 31, 1, 15]. Finally, as datasets for action segmentation are significantly smaller than for other tasks (e.g., action recognition), previous work resorts to data augmentation and self-supervised learning [2, 19, 28]. However, these methods typically augment only local frame features, and do not augment action instances.
@@
To address these limitations, we propose the End-to-End Action Segmentation Transformer (EAST). As shown in Fig. 1, EAST consists of the following modules: (a) a large backbone, (b) a detector that predicts action proposals over coarsely sampled frames; (c) an aggregator that combines the proposals to infer class distributions for all frames at the original (unsampled) frame rate; and (d) a refinement module for predicting the final framewise labels. With EAST, we make the following three key contributions.
First, we enable efficient end-to-end training of EAST by introducing lightweight Contract-Expand Adapters (CEA) into a large backbone network. CEA reduces complexity by compressing and expanding features around depth-wise convolutions. This allows efficient fine-tuning of the large backbone to extract multiscale features directly from raw RGB frames, ensuring they are optimized for action segmentation rather than for unrelated vision tasks. While recent methods such as Bridge-Prompt [20] and FACT [27] also claim end-to-end training from RGB inputs, their training strategy consists of disjoint stages – first training the backbone, then freezing features to train the rest of their segmentation model. In contrast, EAST is trained truly end-to-end, without freezing features at any stage.
Second, unlike most recent work that focuses on framewise labeling without explicitly modeling the temporal extent of action instances, EAST performs action segmentation by detection. Specifically, EAST detects action proposals over a downsampled input by regressing action boundaries over the sampled frames. The proposals are then integrated and refined for the final segmentation. This provides two key advantages: improved efficiency by detecting action proposals on coarsely sampled frames (e.g., 1–6 fps), and enhanced framewise classification by explicitly considering the context of detected action segments. It is worth noting that temporal downsampling does not affect the ground truth or evaluation, as our boundary regression is specified relative to timestamps within the video. These boundary detections are mapped to the original full frame rate, and serve as useful constraints for final action segmentation over all frames.
Third, we introduce a novel proposal-based data augmentation method to enhance EAST training. In training, our experiments show that the high-confidence action proposals generally align well with the ground truth, creating an easier learning signal for the subsequent proposal aggregation and refinement stages. To encourage learning from more challenging cases, we augment training by selectively removing high-confidence proposals before passing the remaining proposals to the subsequent EAST modules.
@@ 3.1 Contract-Expand Adapter
$\bm{\bar{x}}$ $=\bm{\sigma}({\bm{W}_{\rm down}^{\rm\top}}\cdot\bm{x})$ (1)
$\bm{\bar{x}}_{c}$ $=\text{contract}_{HW}(\bm{\bar{x}})$
$\bm{\hat{x}}_{m}$ $={\bm{W}_{\rm mid}^{\rm\top}}\cdot\operatorname{\textbf{DWConv}}_{k}(\bm{\bar{x}}_{c})$
$\bm{\bar{x}}_{e}$ $=\text{expand}_{HW}(\bm{\hat{x}}_{m})$
$\bm{\hat{x}}$ $=\bm{\bar{x}}_{e}+\bm{\bar{x}}$
$\bm{x^{\prime}}$ $=\alpha\cdot{\bm{W}_{\rm up}^{\rm\top}}\cdot\bm{\hat{x}}+\bm{x}$
@@ 3.2 Low-Frame-Rate Action Detection
To enable efficient end-to-end training with temporal downsampling, we adopt a segmentation-by-detection framework, departing from SOTA approaches. Action proposals are predicted on coarsely sampled frames and then integrated at the original high frame rate for framewise classification. Inspired by anchor-free detectors (e.g., FCOS [34] and ActionFormer [41]), we treat each sampled frame as a query for its corresponding action proposal.
This generates high-quality action proposals from low-FPS input. Compared to SOTA methods that rely on separate frame and action branches [27] or learnable queries [4], our approach significantly simplifies training by directly predicting action instances from sampled frames.
@@ 3.4 Training Loss Functions
where $\mathcal{L}_{\text{c}}$ denotes the focal loss [24]; $y_{i}$ is the ground-truth class; $\lambda_{\text{r}}$ is a weighting hyperparameter; $\mathbbm{1}(\cdot)$ is a binary indicator that equals 0 if $i$th frame is classified as background, and 1, otherwise; $T_{+}$ is the total number of sampled frames classified as an action $T_{+}=\sum_{i=1}^{T^{\prime}}\mathbbm{1}(\hat{y}_{i})$; $\mathcal{L}_{\text{r}}$ is the DIoU loss [42] for regression; $t_{n(i)}^{s},t_{n(i)}^{e}]$ denotes the time interval of $n$th ground-truth action instance closest to the predicted interval $\hat{t}_{i}^{s},\hat{t}_{i}^{e}]$ of $i$th action proposal. Note that the regression loss is applied only to action proposals classified as an action excluding background.
@@ 4 Results
[…]
@@ 4.1 Features and Training Efficiency
The SOTA methods use precomputed I3D features [6] for Breakfast and TSM features [23] for Assembly101. To ensure consistency in the frame features across the SOTA methods and EAST, we extracted MAEv2 frame features for all the methods, including ours, using the ViT-G backbone pretrained with VideoMAEv2 [36]. Tab. 1 shows that the SOTA methods improve performance on Breakfast when using precomputed MAEv2 features, while LTContext with MAEv2 features fails to do so on Assembly101. On both datasets, end-to-end trained EAST using the same pretrained ViT-G achieves superior performance.
@@
(’ICCV23) MAEv2* 80.6 75.7 64.1 75.2 76.6
FACT [27]
(’CVPR24)
I3D 81.4 76.5 66.2 79.7 76.2
I3D* 78.4 73.3 62.2 77.4 75.5
MAEv2* 80.6 75.9 65.3 78.9 77.2
EAST MAEv2 85.6 81.5 71.6 83.5 82.2
Table 1: Impact of feature representation on Breakfast. *: results from official code. SOTA methods use precomputed I3D features on Breakfast. To ensure a consistent feature representation, we compare SOTA and EAST using MAEv2 frame features extracted by the ViT-G backbone pretrained with VideoMAEv2 [36] across all the methods. Results are averaged across four standard splits.
@@
Methods MSTCN++ ASFormer LTContext DiffAct FACT EAST
GFLOPS 4.5 7.6 7.8 63.0 5.5 9.2
Table 3: Computational cost in GFLOPS, measured on a 5-minute video using the setup from [27].
Tab. 4 compares per-epoch time, total training time, and GPU memory usage of SOTA methods and EAST on Breakfast using NVIDIA H100s. Note that the SOTA methods in Tab. 4 rely on pre-computed features and in their reports exclude feature extraction time and memory usage. Therefore, the comparison of training efficiency is unfavorable to EAST, whose reported numbers in Tab. 4 account for the full pipeline, including training the large backbone. Tab. 4 shows EAST achieves total training time comparable to LTContext – the fastest feature-based Transformer – despite processing RGB frames.
[…]
@@ 4.2 CEA Adapter and Backbones
Dataset Method F1@10 F1@25 F1@50 Edit Acc
GTEA Baseline 87.3 84.7 75.9 82.5 86.3
MSTCN 93.2 93.0 88.7 92.1 83.8
ASFormer 86.8 84.0 74.3 83.1 83.6
Proposal Aug 95.8 95.4 91.7 95.4 87.1
Breakfast Baseline 78.2 74.2 65.8 75.1 77.6
MSTCN 85.6 81.3 71.5 83.4 80.3
ASFormer 84.1 79.9 69.7 82.2 80.2
Proposal Aug 86.2 82.2 71.8 84.5 82.8
Table 8: Action segmentation results using identical action proposals from our detector, comparing refinement strategies. “Baseline” denotes framewise classification via Eq. 2 without augmentation (Sec. 3.5). “MSTCN” and “ASFormer” apply their respective refinement networks to our generated proposals. “Proposal Aug” (EAST) incorporates our augmentation during training for improved refinement learning. The comparison between Baseline and Proposal Aug isolates our proposed augmentation’ impact.
@@ 4.4 Comparison with SOTA
ViT-G fine-tuned EAST 93.1 91.9 88.6 88.4 91.9
@@
ViT-G fine-tuned EAST 86.2 82.2 71.8 84.5 82.8
@@ 4.5 Qualitative Results
Fig. 4 compares of EAST segmentation results with the ground truth and SOTA on sample videos from GTEA, Breakfast, and Assembly101. In the top example video from GTEA, EAST successfully detects all actions, whereas the other methods miss at least one action. On the middle example video from Breakfast, the SOTA methods exhibit varying degrees of missed actions, spurious predictions, or oversegmentation. In contrast, EAST provides high-quality segmentation closely aligned with the ground truth. Finally, in the bottom video from Assembly101, EAST fails to detect a brief action instance and incorrectly identifies an action boundary too early – both of these are ambiguous edge cases that are challenging to discern even through visual inspection.
Figure 4: Segmentation results on sample videos from GTEA (top), Breakfast (middle), and Assembly101 (bottom). For each video sequence, the top row shows the color-coded ground-truth action instances, middle row shows EAST output, and bottom row shows SOTA results generated using publicly available models.
@@ 5 Conclusion
We have introduced EAST – the first fully end-to-end trainable action segmenter. EAST performs segmentation by detection, which enables temporal downsampling of input videos, significantly reducing computational costs. EAST takes RGB frames, sampled at a low frame rate, as input to a large-scale backbone. The backbone is fine-tuned using our Contract-Expand Adapter (CAE). CAE is especially effective in reducing the computational costs of end-to-end training by leveraging spatial pooling. The backbone features are passed to the detector to predict action proposals, which are then aggregated and refined to produce final framewise labeling at the original unsampled frame rate. We have also specified a novel proposal-based data augmentation that increases uncertainty of the detector’s output during training, effectively simulating test-time conditions. EAST outperforms prior work across all metrics on the GTEA, 50Salads, Breakfast, and Assembly101 datasets, while maintaining comparable processing times, even with the additional step of extracting frame features by the backbone. We have conducted a comprehensive ablation study to evaluate EAST’s performance under various configurations. Additional results are presented in the supplements.
Acknowledgement: This work has been supported by USDA NIFA award No.2021-67021-35344.
@@ 6 Additional Results
SOTA Methods with Low FPS Input. The impact of video downsampling on the SOTA methods and EAST is evaluated on Breakfast in Table 15. The SOTA methods are trained using I3D frame features sampled at 1 FPS, with their output framewise classification subsequently upscaled to the original 15 FPS, as per their reported evaluation setting. Table 15 shows significant performance degradation for all the SOTA methods when working with the low FPS at the input. As shown in Tables 15 and 9, EAST maintains the best performance at low FPS rates and improves as the frame rate increases, subject to memory and compute constraints.